OpenStudy (tay0913):

2(8-x)=2 and there's more😕

OpenStudy (anonymous):

you can start by dividing both sides by 2.

OpenStudy (cowgirl1):

ok sirt you distribute the the left side ie. 3(3-2) = 9-36

OpenStudy (cowgirl1):

er 9-6

alones (alones):

Hy welcoe to Openstudy oka so actually you can use this site it will provide staps and answer :) http://mathpapa.com/algebra-calculator.html

OpenStudy (tay0913):

I didn't understand u cowgirl1

OpenStudy (cowgirl1):

ok so like 2(8-x)=2, right? that = (2x8)-2x=2

OpenStudy (cowgirl1):

so 2(8-x)=2 becomes 16-2x=2

OpenStudy (tay0913):

ok.

OpenStudy (cowgirl1):

so next we can bring x to one side by adding 2x to each side. do you know that part?

OpenStudy (cowgirl1):

16-2x=2 will become 16=2+2x

OpenStudy (tay0913):

Mhm

OpenStudy (cowgirl1):

subtract 2 from each side 16=2+2x will be 14=2x

OpenStudy (cowgirl1):

divide each side by 2 14=2x will be 7=x

OpenStudy (tay0913):

Thx

OpenStudy (cowgirl1):

youre welcome :-)

rebeccaxhawaii (rebeccaxhawaii):

\[\frac{ 2(8-x) }{ 2 }=\frac{ 2 }{ 2 }\] \[2(8-x)=1\] \[8-1=x\] \[x=7\]

rebeccaxhawaii (rebeccaxhawaii):

here this is a little more easy to read

OpenStudy (tay0913):

Thx can u help me with another

OpenStudy (cowgirl1):

yeah

OpenStudy (tay0913):

5x-1=2x+5

rebeccaxhawaii (rebeccaxhawaii):

5x-2x=5+1 3x=6 x=2

OpenStudy (tay0913):

Thx how about 14 more

OpenStudy (cowgirl1):

wow that was fast

rebeccaxhawaii (rebeccaxhawaii):

choose best response then open a new question